OpenStudy (anonymous):

the equation that represents the line passing through (-3,-1) with a slope of 4 is?

OpenStudy (anonymous):

The equation we will want to use is: \[y-y_1=m(x-x_1)\] where m is the slope, and x_1 and y_1 are the coordinates of a point on that line. We have a point and the slope, so we can plug in: \[y+1=4(x+3)\] And after simplifying: \[y=4x+11\]

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

No problem, glad I could help.